OpenStudy (anonymous):

Solve the following system of equations by elimination: 3x + 2y= 8 x + y= 2. A. x= -4 y= -2 B. x= 4 y= -2 C. x= 4 y= 1 D. x= 6 y= 3

OpenStudy (anonymous):

Awarding a medal!!!

OpenStudy (anonymous):

@alexwright can you help me?

OpenStudy (anonymous):

@johnweldon1993 could you help me??

OpenStudy (anonymous):

@helpout3 can you help me?

OpenStudy (anonymous):

No i cant help sorry Math is not my strong suit

OpenStudy (anonymous):

Aah ok, thank you very much!! =)

OpenStudy (texaschic101):

3x + 2y = 8 x + y = 2 ---->(-3) ----------------- 3x + 2y = 8 -3x - 3y = -6 (result of multiplying by -3) -----------------add -y = 2 y = -2 now sub -2 in for y in either equation to find x x + y = 2 x - 2 = 2 x = 4 solution : x = 4 and y = -2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Thank you, =)

OpenStudy (texaschic101):

sure thing :)